Alex has 1400 ft of irrigation pipping. He wants to use it to irrigate his back lawn. He wants to lay the pipping in such a manner as to cut off 3 equal size rectangle regions in the yard. What are the dimensions that would produce the maximum enclosed area.
Answer by TimothyLamb(4379)

---
2x + 2y = 1400
y = (1400 - 2x)/2
y = 700 - x
a = xy
---
a = xy
a = x(700 - x)
a = 700x - xx
a = -xx + 700x
---
the above quadratic equation is in standard form, with a=-1, b=700, and c=0
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
-1 700 0
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic vertex is a maximum at: ( x=350, a= 122500 )
---
answer:
length = 350 ft
width = 350 ft
maximum area = 122500 sq.ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php